OpenStudy (anonymous):

Can somebody check these 5 questions for me?

OpenStudy (anonymous):

I meant 4 questions

OpenStudy (anonymous):

<< = the answer i selected

OpenStudy (anonymous):

Solve the equation. If necessary, round your answer to the nearest tenth. x^2 = 324 A. –18, 18 << B. 18 C. –162, 162 D. 162

OpenStudy (anonymous):

2. Solve the equation. If necessary, round your answer to the nearest tenth. 4x^2 = 121 A. 2.75, –2.75 B. 5.5, –5.5 C. 11, –11 << D. 30.25, –30.25

OpenStudy (anonymous):

Solve the equation. x^2 = 160 A B C: 14 D:-14,14

OpenStudy (anonymous):

I choose at

OpenStudy (anonymous):

What is the value of y? 4y^2 + 8 = 264 ± 8

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (texaschic101):

x^2 = 324 x = sq rt 324 x = -18,18

OpenStudy (texaschic101):

4x^2 = 121 --- divide both sides by 4 x^2 = 30.25 x = sq rt 30.25 = -5.5, 5.5

OpenStudy (texaschic101):

x^2 = 160 x = sq rt 160 x = 4 sq rt 10, -4 sq rt 10

OpenStudy (texaschic101):

4x^2 + 8 = 264 --- subtract 8 from both sides 4x^2 = 256 -- divide both sides by 4 x^2 = 64.....x = sq rt 64.....x = 8,-8

OpenStudy (anonymous):

thank you